# Optimal Non-Adaptive Group Testing with One-Sided Error Guarantees

### v3

@@ III-A Proof of Theorem 2 (Achievability for SUBSET)
Note that when $\theta>\frac{1}{3}$ and $\eta^{-}=\Theta((\frac{k}{n})^{\lambda})$ for some $\lambda\in(0,\frac{1}{2})$ as assumed in the lemma statement, the expression $\eta^{-}k\log\frac{1}{\eta^{-}}$ tends to $\infty$ as $n\to\infty$, meaning that the bound in (23) is $o(1)$. Under these assumptions, we also have $\eta^{-}k\log\frac{1}{\eta^{-}}=\omega(\frac{1}{\sqrt{T}})$ when $T=\big(k\log_{2}\frac{n}{k}\big)(1+\varepsilon)$ (via $k\geq n^{\frac{1}{3}+\Omega(1)}$), so we can take this as the value of $\delta_{2}$ in (i).
@@
We now use (23) along with a union bound to obtain a counterpart that holds simultaneously for all $\bar{s}\in{\mathcal{S}}_{\mathrm{an}}$. The distribution on the number of tests in which $\bar{s}$ is present is the same for any $\bar{s}$, and there are at most ${k\choose(1-\eta^{-})k}={k\choose\eta^{-}k}$ sets in ${\mathcal{S}}_{\mathrm{an}}$ (recall that we implicitly condition on fixed $S=s$), so the probability of any such set being present in less than $T(1-2^{-(1-\eta^{-})(1+o(1))})(1-\delta_{2})$ tests is upper bounded by
${k\choose\eta^{-}k}\exp\left(-2\eta^{-}k\log\frac{1}{\eta^{-}}\right)$ (24)
$=$ $\exp\left(\eta^{-}k\log\frac{1}{\eta^{-}}(1+o(1))-2\eta^{-}k\log\frac{1}{\eta^{-}}\right)$ (25)
$=$ $\exp\left(-\eta^{-}k\log\frac{1}{\eta^{-}}(1-o(1))\right)$ (26)
@@
where (25) follows since $\eta^{-}k=o(k)$ (i.e., $\eta^{-}=o(1)$), and (27) follows from the assumptions $\eta^{-}=\Theta((\frac{k}{n})^{\lambda})$, $\lambda\in\big(0,\frac{1}{2}\big)$, and $\theta>\frac{1}{3}$. Thus, all sets in ${\mathcal{S}}_{\mathrm{an}}$ are present in at least $T(1-2^{-(1-\eta^{-})(1+o(1))})(1-\delta_{2})$ tests. ∎
Lemma 6 hinges on the two specified assumptions on $\eta^{-}$ and $\theta$; we now proceed to argue that it is safe to assume both of these for the purpose of proving Theorem 2. For $\eta^{-}$, this is because we only replaced the assumed scaling from $\Omega(\cdot)$ to $\Theta(\cdot)$ (and restricted $\lambda$ to $(0,\frac{1}{2})$ rather than $0,\frac{1}{2})$); however, solving SUBSET with a smaller $\eta^{-}$ immediately implies the same for larger $\eta^{-}$, since one can simply remove an arbitrary subset of items from the final estimate. For $\theta$, the restriction $\theta>\frac{1}{3}$ is valid because when $\theta\leq\frac{1}{3}$, a rate of $1$ is achievable even for exact recovery [18, 16], which implies solving SUBSET for arbitrary $\eta^{-}$. Hence, for the rest of the proof, we adopt the stricter assumptions from Lemma 6, in particular noting that $\eta^{-}=o(1)$.
Using part (ii) of Lemma 6 and the fact that $\tilde{S}\in{\mathcal{S}}_{\mathrm{an}}$, we have that $\tilde{S}$ explains at least $T(1-2^{-(1-\eta^{-})(1+o(1))})$ tests with probability 1o(1)1-o(1) (since $\tilde{S}$ explains every test it is present in due to $\tilde{S}\subseteq S$). Hence, and using part (i) of Lemma 6, the number of positive tests not explained by $\tilde{S}$ is at most the following, with probability 1o(1)1-o(1):
$\frac{T}{2}(1+\delta_{2})-T(1-2^{-(1-\eta^{-})(1+o(1))})(1-\delta_{2})=\frac{T}{2}\big(2^{\eta^{-}(1+o(1))}-1+\delta_{2}(3-2^{\eta^{-}(1+o(1))})\big).$ (28)
@@
$\delta_{2}=\sqrt{\frac{16\eta^{-}k\log\frac{1}{\eta^{-}}}{T}}=\Theta\bigg(\sqrt{\frac{\eta^{-}\log\frac{1}{\eta^{-}}}{\log n}}\bigg)=\Theta\big(\sqrt{\eta^{-}}\big)=n^{-\Omega(1)},$ (29)
where we applied $T=\Theta(k\log n)$ and $\log\frac{1}{\eta^{-}}=\Theta(\log n)$ (via $\eta^{-}=\Theta((\frac{k}{n})^{\lambda})$).
@@
Our proof strategy consists of finding a “worst case distribution” on the number of positive tests the non-defective items in $\mathcal{N}$$_{\ell}$ get placed in, and showing that the probability of $\tilde{S}^{\prime}$ explaining more tests than $\tilde{S}$ under this distribution is small. The probability of at least one of the $\ell$ non-defective items in $\mathcal{N}$$_{\ell}$ being included in any given test is $1-\big(1-\frac{\log 2(1+o(1))}{k}\big)^{\ell}=\frac{\ell\log 2}{k}\big(1+o(1)\big)$, since $\ell=o(k)$. Thus, recalling the high probability upper bound on the number of tests left unexplained by $\tilde{S}$ given in (28), the number of newly explained tests is stochastically upper bounded by the following: [^3] Our analysis focuses on the positive tests that these $\ell$ non-defectives get placed in. Any negative tests they get placed in will only decrease the number of explained tests further, since by Definition 4 any non-defective in a negative test does not explain any tests. However, for our purposes it suffices to naively upper bound this amount of decrease by zero.
$E\sim\mathrm{Bin}\bigg(\frac{T}{2}\big(2^{\eta^{-}(1+o(1))}-1+\delta_{2}(3-2^{\eta^{-}(1+o(1))})\big),\frac{\ell\log 2}{k}(1+o(1))\bigg).$ (31)
Recalling the second part of Lemma 5, we wish to analyze the probability that $E\geq\frac{1-\delta_{1}}{2}\ell\log\frac{n}{k}$. To do so, we first analyze the mean
$\mu_{E}\coloneqq\mathbb{E}[E]=\frac{T\big(2^{\eta^{-}(1+o(1))}-1+\delta_{2}(3-2^{\eta^{-}(1+o(1))})\big)\ell\log 2}{2k}\big(1+o(1)\big).$ (32)
We also note the following asymptotic simplification for $2^{\eta^{-}(1+o(1))}-1$ when $\eta^{-}=o(1)$:
$2^{\eta^{-}(1+o(1))}$ $=e^{\eta^{-}\log 2(1+o(1))}=1+\eta^{-}\log 2\big(1+o(1)\big)+O\left((\eta^{-})^{2}\right)$
$\implies 2^{\eta^{-}}-1=\eta^{-}\log 2\big(1+o(1)\big).$ (33)
@@
As a result, the term $2^{\eta^{-}}-1=O(\eta^{-})$ is dominated by the expression $\delta_{2}(3-2^{\eta^{-}})=\Theta(\delta_{2})=\Theta\big(\sqrt{\eta^{-}})$ (see (29)). Thus, we have the following asymptotic expression for the mean $\mu_{E}$:
$\mu_{E}$ $=\frac{T\delta_{2}(3-2^{\eta^{-}(1+o(1))})\ell\log 2}{2k}\big(1+o(1)\big)$ (34)
$=\frac{\left(k\log_{2}\frac{n}{k}\right)(1+\varepsilon)\ell\delta_{2}(3-2^{\eta^{-}(1+o(1))})\log 2}{2k}\big(1+o(1)\big)$ (35)
$=\frac{(\log\frac{n}{k})(1+\varepsilon)\ell\delta_{2}(3-2^{\eta^{-}(1+o(1))})}{2}\big(1+o(1)\big),$ (36)
@@
where (35) follows by substituting $T=\big(k\log_{2}\frac{n}{k}\big)(1+\varepsilon)$. Hence, the scaling of $\mu_{E}$ is $\Theta(\ell\delta_{2}\log n)$. Comparing this to the scaling of the minimum number of tests that needs to be explained, which is $\frac{1-\delta_{1}}{2}\ell\log\frac{n}{k}=\Theta((1-\delta_{1})\ell\log n)$ (recall that Lemma 4 holds for any $\delta_{1}=1-o(1)$). Choosing $\delta_{1}=1-\frac{1}{\sqrt{\log n}}$, this scaling thus becomes $\Theta(\ell\sqrt{\log n})$. Thus, we see that $\mu_{E}$ scales more slowly, since $\delta_{2}=n^{-\Omega(1)}$) (see (29)). As such, we can use binomial concentration bounds to show that $\mathbb{P}(E\geq\frac{1-\delta_{1}}{2}\ell\log\frac{n}{k})$ is small.
@@
To be more specific, we write $\frac{1-\delta_{1}}{2}\ell\log\frac{n}{k}=(1+\delta_{3})\mu_{E}$ for some $\delta_{3}>0$. Due to the aforementioned scaling laws on $1-\delta_{1}$ and $\delta_{2}$, we have that $\delta_{3}=\Theta(\frac{1-\delta_{1}}{\delta_{2}})=\frac{n^{\Omega(1)}}{\sqrt{\log n}}=\omega(1)$. Thus, by the multiplicative Chernoff bound (Appendix -A), we have
$\mathbb{P}\left(E\geq\frac{1-\delta_{1}}{2}\ell\log\frac{n}{k}\right)$ $=\mathbb{P}(E\geq(1+\delta_{3})\mu_{E})$ (37)
$\leq\exp\left(-\big((1+\delta_{3})\log(1+\delta_{3})-\delta_{3}\big)\mu_{E}\right)$ (38)
$=\exp\left(-(1+\delta_{3})\log(1+\delta_{3})\mu_{E}\big(1-o(1)\big)\right)$ (39)
$=\exp\left(-\Omega\Big(\ell\sqrt{\log n}\cdot\log\delta_{3}\Big)\right)$ (40)
$=\exp\left(-\Omega\left(\ell\sqrt{\log n}\cdot\log n\right)\right),$ (41)
@@
where (39) follows since $\delta_{3}=\omega(1)$, (40) follows since $(1+\delta_{3})\mu_{E}=\frac{1-\delta_{1}}{2}\ell\log n=\Omega(\ell\sqrt{\log n})$, and (41) follows since $\delta_{3}=\frac{n^{\Omega(1)}}{\sqrt{\log n}}$ and hence $\log\delta_{3}=\Omega(\log n)$. Lemma 7 then follows by recalling that $\mathbb{P}(T_{\exp}(\tilde{S})\leq T_{\exp}(\tilde{S}^{\prime}))\leq\mathbb{P}(E\geq\frac{1-\delta_{1}}{2}\ell\log n)$ (see Lemma 5). ∎
@@
With this result established, we are now in a position to show that a union bound over all possible $({\mathcal{D}}_{\ell},{\mathcal{N}}_{\ell})$, followed by a second union bound over all the values of $\ell$, still leads to the probability of $\tilde{S}$ not explaining the most tests being small. This is formalized in the following lemma.
@@
Under the setup and notation of Lemma 7, [^4] In particular, we are again conditioning on (21) and (28), and the probability is over the randomness in the placements of non-defective items. for each $\ell\in\{1,2,\dots,5\eta^{-}k\}$, define
${\mathcal{E}}_{\ell}\coloneqq\bigcup_{{\mathcal{D}}_{\ell}\subset\tilde{S}\,:\,\lvert{\mathcal{D}}_{\ell}\rvert=\ell}\bigcup_{{\mathcal{N}}_{\ell}\subset S^{\mathsf{c}}\,:\,\lvert{\mathcal{N}}_{\ell}\rvert=\ell}\left\{T_{\exp}(\tilde{S})\leq T_{\exp}((\tilde{S}\setminus{\mathcal{D}}_{\ell})\cup{\mathcal{N}}_{\ell})\right\}$ (42)
as the event that at least one way of swapping out $\ell$ items from $\tilde{S}$ with $\ell$ non-defective items leads to more tests being explained. Then, we have the following as $n\to\infty$:
$\mathbb{P}\left(\bigcup_{\ell=1}^{5\eta^{-}k}{\mathcal{E}}_{\ell}\right)\to 0.$ (43)
@@
We begin by deriving a bound on $\mathbb{P}({\mathcal{E}}_{\ell})$ for fixed $\ell$. For the event corresponding to any single choice of $({\mathcal{D}}_{\ell},{\mathcal{N}}_{\ell})$ in (42), the associated probability is upper bounded via Lemma 7. A simple counting argument reveals that there are ${(1-\eta^{-})k\choose\ell}{n-k\choose\ell}$ such pairs, which can be further upper bounded by ${k\choose\ell}{n\choose\ell}$. Hence, the union bound gives
$\mathbb{P}({\mathcal{E}}_{\ell})$ $\leq{k\choose\ell}{n\choose\ell}\mathbb{P}\left(E\geq\frac{1-\delta_{1}}{2}\ell\log\frac{n}{k}\right)$ (44)
$\leq{k\choose\ell}{n\choose\ell}\exp\left(-\Omega\big(\ell\sqrt{\log n}\cdot\log n\big)\right)$ (45)

### v4

@@ III-A Proof of Theorem 2 (Achievability for SUBSET)
Note that when $\theta>\frac{1}{3}$ and $\eta=\Theta((\frac{k}{n})^{\lambda})$ for some $\lambda\in(0,\frac{1}{2})$ as assumed in the lemma statement, the expression $\eta k\log\frac{1}{\eta}$ tends to $\infty$ as $n\to\infty$, meaning that the bound in (28) is $o(1)$. Consequently, we also have $\sqrt{\frac{16(3+\frac{1}{\lambda})\eta k\log\frac{1}{\eta}}{T}}=\omega(\frac{1}{\sqrt{T}})$, so we can take this as the value of $\delta_{2}$ in (i).
We now use (28) along with a union bound to obtain a counterpart that holds simultaneously for all $\bar{s}\in{\mathcal{S}}_{\mathrm{an}}$. To do so, we first require an upper bound on the size of ${\mathcal{S}}_{{\rm an}}$. Toward this, observe that we can calculate $\lvert{\mathcal{S}}_{{\rm an}}\rvert$ directly via the following counting argument:
  • Consider $\bar{s}\in{\mathcal{S}}_{{\rm an}}$ as being formed by taking $(1-\eta)k-i$ items from $S$ (recall that we implicitly condition on fixed $S=s$) and $i$ items from $S^{\mathsf{c}}$ for some $i\geq 0$.
  • Since $\lvert\bar{s}\rvert=(1-\eta)k$, the Hamming distance between $\bar{s}$ and $S$ satisfies $d_{\mathrm{H}}(\bar{s},S)=\eta k+2i$. Moreover, since $\bar{s}\in{\mathcal{S}}_{{\rm an}}$, it satisfies $d_{\mathrm{H}}(\bar{s},S)\leq 5\eta k$. Thus, it holds that $0\leq i\leq 2\eta k$.
  • For each $i$, the number of possible $\bar{s}\in{\mathcal{S}}_{{\rm an}}$ that can be formed this way is ${k\choose(1-\eta)k-i}{n-k\choose i}={k\choose\eta k+i}{n-k\choose i}$.
  • Collectively this construction for $i=0,\dots,2\eta k$ produces all sets in ${\mathcal{S}}_{{\rm an}}$, and it follows that
    $\lvert{\mathcal{S}}_{{\rm an}}\rvert=\sum_{i=0}^{2\eta k}{k\choose\eta k+i}{n-k\choose i}$. (29)
We now proceed to bound $\lvert{\mathcal{S}}_{{\rm an}}\rvert$:
$\lvert{\mathcal{S}}_{{\rm an}}\rvert$ $=\sum_{i=0}^{2\eta k}{k\choose\eta k+i}{n-k\choose i}$ (30)
$\leq(2\eta k+1){k\choose 3\eta k}{n\choose 2\eta k}$ (31)
$=(2\eta k+1)\exp\Big(3\eta k\log\Big(\frac{1}{3\eta}\Big)(1+o(1))$
$\qquad\qquad+2\eta k\log\Big(\frac{n}{2\eta k}\Big)(1+o(1))\Big)$ (32)
$=\exp\left(\Big(5+\frac{2}{\lambda}\Big)\eta k\log\Big(\frac{1}{\eta}\Big)(1+o(1))\right),$ (33)
where (31) follows from bounding $n-k\choose i$} by $n\choose i$} and since the largest term in the sum is when $i=2\eta k$, (32) holds since $\eta k=o(k)$ (i.e., $\eta=o(1)$), and (33) follows from writing $\frac{n}{k}=\Theta((\frac{1}{\eta})^{\frac{1}{\lambda}})$ and simplifying, as well as absorbing the prefactor $2\eta k+1$ into the $o(1)$ term.
@@
With the bound on $\lvert{\mathcal{S}}_{{\rm an}}\rvert$ in place, we are now in a position to take a union bound over all $\bar{s}\in{\mathcal{S}}_{{\rm an}}$. The distribution of the number of tests in which $\bar{s}$ is present is the same for each $\bar{s}$ and satisfies the tail bound in (28), and the number of these sets is upper bounded via (33). Thus, the probability of any such set being present in less than $T(1-2^{-(1-\eta)(1+o(1))})(1-\delta_{2})$ tests is upper bounded by
$\exp\left(\Big(5+\frac{2}{\lambda}\Big)\eta k\log\Big(\frac{1}{\eta}\Big)(1+o(1))\right)$
$\qquad\qquad\qquad\times\exp\left(-2\Big(3+\frac{1}{\lambda}\Big)\eta k\log\frac{1}{\eta}\right)$ (34)
$=\exp\left(-\eta k\log\Big(\frac{1}{\eta}\Big)(1-o(1))\right)$ (35)
@@
where (36) follows since $\eta k\log\frac{1}{\eta}\to\infty$ as established following (28). Thus, all sets in ${\mathcal{S}}_{\mathrm{an}}$ are present in at least $T(1-2^{-(1-\eta)(1+o(1))})(1-\delta_{2})$ tests. ∎
Lemma 6 hinges on the two specified assumptions on $\eta$ and $\theta$; we now proceed to argue that it is safe to assume both of these for the purpose of proving Theorem 2. For $\eta$, this is because we only replaced the assumed scaling from $\Omega(\cdot)$ to $\Theta(\cdot)$ (and restricted $\lambda$ to $(0,\frac{1}{2})$ rather than $0,\frac{1}{2})$); however, solving SUBSET with a smaller $\eta$ immediately implies the same for larger $\eta$, since one can simply remove an arbitrary subset of items from the final estimate. For $\theta$, the restriction $\theta>\frac{1}{3}$ is valid because when $\theta\leq\frac{1}{3}$, a rate of $1$ is achievable even for exact recovery [28, 10], which implies solving SUBSET for arbitrary $\eta$. Hence, for the rest of the proof, we adopt the stricter assumptions from Lemma 6, in particular noting that $\eta=o(1)$.
Using part (ii) of Lemma 6 and the fact that $\tilde{S}\in{\mathcal{S}}_{\mathrm{an}}$, we have that $\tilde{S}$ explains at least $T(1-2^{-(1-\eta)(1+o(1))})$ tests with probability 1o(1)1-o(1) (since $\tilde{S}$ explains every test it is present in due to $\tilde{S}\subseteq S$). Hence, and using part (i) of Lemma 6, the number of positive tests not explained by $\tilde{S}$ is at most the following, with probability 1o(1)1-o(1):
$\frac{T}{2}(1+\delta_{2})-T(1-2^{-(1-\eta)(1+o(1))})(1-\delta_{2})$
$=\frac{T}{2}\big(2^{\eta(1+o(1))}-1+\delta_{2}(3-2^{\eta(1+o(1))})\big).$ (37)
@@
$\delta_{2}=\sqrt{\frac{16{(3+\frac{1}{\lambda})}\eta k\log\frac{1}{\eta}}{T}}=\Theta\bigg(\sqrt{\frac{\eta\log\frac{1}{\eta}}{\log n}}\bigg)$ $=\Theta\big(\sqrt{\eta}\big)$
$=n^{-\Omega(1)},$ (38)
where we applied $T=\Theta(k\log n)$ and $\log\frac{1}{\eta}=\Theta(\log n)$ (via $\eta=\Theta((\frac{k}{n})^{\lambda})$).
@@
Our proof strategy consists of finding a “worst case distribution” on the number of positive tests the non-defective items in $\mathcal{N}$$_{\ell}$ get placed in, and showing that the probability of $\tilde{S}^{\prime}$ explaining more tests than $\tilde{S}$ under this distribution is small. The probability of at least one of the $\ell$ non-defective items in $\mathcal{N}$$_{\ell}$ being included in any given test is $1-\big(1-\frac{\log 2(1+o(1))}{k}\big)^{\ell}=\frac{\ell\log 2}{k}\big(1+o(1)\big)$, since $\ell=o(k)$. Thus, recalling the high probability upper bound on the number of tests left unexplained by $\tilde{S}$ given in (37), we have that the number of tests that could be newly explained by $\mathcal{N}$$_{\ell}$ is at most
$\frac{T}{2}\big(2^{\eta(1+o(1))}-1+\delta_{2}(3-2^{\eta(1+o(1))})\big)+\Psi_{{\mathcal{D}}_{\ell}},$ (40)
where $\Psi_{{\mathcal{D}}_{\ell}}\coloneqq T_{{\rm exp}}(\tilde{S})-T_{{\rm exp}}(\tilde{S}\setminus{\mathcal{D}}_{\ell})$ is the number of previously-explained tests that are no longer explained after removing the items in $\mathcal{D}$$_{\ell}$ from $\tilde{S}$. We require the inclusion of $\Psi_{{\mathcal{D}}_{\ell}}$ in (40) since it concerns tests no longer explained following the removal of items in $\mathcal{D}$$_{\ell}$, and these tests may be “re-explained” by an item in $\mathcal{N}$$_{\ell}$. Also note that since we are conditioned on (26), Lemma 5 implies that $\Psi_{{\mathcal{D}}_{\ell}}\geq\frac{1-\delta_{1}}{2}\ell\log\frac{n}{k}$ for some $\delta_{1}=1-o(1)$ (to be specified later), which will be useful later on in the proof.
From (40), it follows that the number tests that are newly explained by $\mathcal{N}$$_{\ell}$ (including those “re-explained”) is stochastically dominated [^3] For two random variables $X_{1},X_{2}$, we say that $X_{1}$ is stochastically dominated by $X_{2}$ (or $X_{2}$ stochastically dominates $X_{1}$) if $\mathbb{P}(X_{1}\geq x)\leq\mathbb{P}(X_{2}\geq x),\ \forall x\in\mathbb{R}$. This means that $X_{2}$ is “greater than or equal to $X_{1}$” in a probabilistic sense. by the following: [^4] Our analysis focuses on the positive tests that these $\ell$ non-defectives get placed in. Any negative tests they get placed in will only decrease the number of explained tests further, since by Definition 4 any non-defective in a negative test does not explain any tests. However, for our purposes it suffices to use a naive upper bound that replaces this decrease by zero.
$E\sim\mathrm{Bin}\bigg(\frac{T}{2}\big(2^{\eta(1+o(1))}-1+\delta_{2}(3-2^{\eta(1+o(1))})\big)$
$+\Psi_{{\mathcal{D}}_{\ell}},\frac{\ell\log 2}{k}(1+o(1))\bigg).$ (41)
Moreover, the definitions of $E$ and $\Psi_{{\mathcal{D}}_{\ell}}$ directly imply that
$\{T_{\exp}(\tilde{S})\leq T_{\exp}(\tilde{S}^{\prime})\}=\{E\geq\Psi_{{\mathcal{D}}_{\ell}}\},$ (42)
since when forming $\tilde{S}^{\prime}$ from $\tilde{S}$ there is a “loss” of $\Psi_{{\mathcal{D}}_{\ell}}$ from removing $\mathcal{D}$$_{\ell}$, but then a “gain” of $E$ from adding $\mathcal{N}_{\ell}$.
Thus, it suffices to upper bound $\mathbb{P}(E\geq\Psi_{{\mathcal{D}}_{\ell}})$. To do so, we first analyze the mean:
$\mu_{E}$ $\coloneqq\mathbb{E}[E]$
$=\Bigg(\frac{T\big(2^{\eta(1+o(1))}-1+\delta_{2}(3-2^{\eta(1+o(1))})\big)\ell\log 2}{2k}$
$\qquad\qquad\qquad\qquad+\frac{\ell\Psi_{{\mathcal{D}}_{\ell}}\log 2}{k}\Bigg)\big(1+o(1)\big).$ (43)
We also note the following asymptotic simplification for $2^{\eta(1+o(1))}-1$ when $\eta=o(1)$:
$2^{\eta(1+o(1))}$ $=e^{\eta\log 2(1+o(1))}=1+\eta\log 2\big(1+o(1)\big)$
$\implies 2^{\eta(1+o(1))}-1=\eta\log 2\big(1+o(1)\big).$ (44)
@@
As a result, the term $2^{\eta}-1=O(\eta)$ is dominated by the expression $\delta_{2}(3-2^{\eta})=\Theta(\delta_{2})=\Theta\big(\sqrt{\eta})$ (see (38)). Thus, we have the following asymptotic expression for the mean $\mu_{E}$:
$\mu_{E}$ $=\bigg(\frac{T\delta_{2}(3-2^{\eta(1+o(1))})\ell\log 2}{2k}$
$\qquad\qquad\qquad\qquad\quad+\frac{\ell\Psi_{{\mathcal{D}}_{\ell}}\log 2}{k}\bigg)\big(1+o(1)\big)$ (45)
$=\bigg(\frac{\left(k\log_{2}\frac{n}{k}\right)(1+\varepsilon)\ell\delta_{2}(3-2^{\eta(1+o(1))})\log 2}{2k}$
$\qquad\qquad\qquad\qquad\quad+\frac{\ell\Psi_{{\mathcal{D}}_{\ell}}\log 2}{k}\bigg)\big(1+o(1)\big)$ (46)
$=\bigg(\frac{(\log\frac{n}{k})(1+\varepsilon)\ell\delta_{2}(3-2^{\eta(1+o(1))})}{2}$
$\qquad\qquad\qquad\qquad\quad+\frac{\ell\Psi_{{\mathcal{D}}_{\ell}}\log 2}{k}\bigg)\big(1+o(1)\big),$ (47)
where (46) follows by substituting $T=\big(k\log_{2}\frac{n}{k}\big)(1+\varepsilon)$. Hence, the scaling of $\mu_{E}$ is
$\mu_{E}=\Theta\left(\ell\Big(\delta_{2}\log n+\frac{\Psi_{{\mathcal{D}}_{\ell}}}{k}\Big)\right).$ (48)
We can compare this to the above-established threshold of $\Psi_{{\mathcal{D}}_{\ell}}$. Defining $\delta_{3}>0$ to satisfy $\Psi_{{\mathcal{D}}_{\ell}}=(1+\delta_{3})\mu_{E}$, we have via (48) that
$\Psi_{{\mathcal{D}}_{\ell}}$ $=(1+\delta_{3})\mu_{E}$ (49)
$=\Theta\left((1+\delta_{3})\ell\Big(\delta_{2}\log n+\frac{\Psi_{{\mathcal{D}}_{\ell}}}{k}\Big)\right)$ (50)
$\implies 1+\delta_{3}=\Theta\left(\frac{\Psi_{{\mathcal{D}}_{\ell}}}{\ell(\delta_{2}\log n+\frac{\Psi_{{\mathcal{D}}_{\ell}}}{k})}\right).$ (51)
Since the quantity in the right hand side of (51) is increasing in $\Psi_{{\mathcal{D}}_{\ell}}$, and recalling that $\Psi_{{\mathcal{D}}_{\ell}}\geq\frac{1-\delta_{1}}{2}\ell\log\frac{n}{k}$ when conditioned on (26), we can therefore deduce that
$1+\delta_{3}$ $=\Omega\left(\frac{\frac{1-\delta_{1}}{2}\ell\log\frac{n}{k}}{\ell(\delta_{2}\log n+\frac{(1-\delta_{1})\ell\log\frac{n}{k}}{2k})}\right)$ (52)
$=\Omega\left(\frac{1}{\frac{\delta_{2}}{1-\delta_{1}}+\frac{\ell}{k}}\right)$ (53)
$=\Omega\left(\frac{1}{\frac{\delta_{2}}{1-\delta_{1}}+\eta}\right)$ (54)
$=\Omega\left(\frac{1-\delta_{1}}{\delta_{2}}\right),$ (55)
@@
where (53) is obtained by dividing through by $\frac{1-\delta_{1}}{2}\ell\log\frac{n}{k}$ and simplifying, (54) uses $\ell=O(\eta k)$ and the fact that the fraction is decreasing in $\ell$, and (55) follows since $\delta_{2}=\Theta(\sqrt{\eta})=\omega(\eta)$ (see (38)). Choosing $\delta_{1}=1-\frac{1}{\sqrt{\log n}}$ and using the scaling of $\delta_{2}$ in (38), the scaling in (55) becomes $1+\delta_{3}=\Omega(\frac{1}{\delta_{2}\sqrt{\log n}})=\frac{n^{\Omega(1)}}{\sqrt{\log n}}=\omega(1)$, which implies the same scaling for $\delta_{3}$. Thus, we see that $\mu_{E}$ grows much more slowly than $\Psi_{{\mathcal{D}}_{\ell}}$. As such, we can use binomial concentration bounds to show that $\mathbb{P}(E\geq\Psi_{{\mathcal{D}}_{\ell}})$ is small.
@@
To be more specific, we use the aforementioned scaling of $\delta_{3}$, the fact that $\Psi_{{\mathcal{D}}_{\ell}}=(1+\delta_{3})\mu_{E}$, and the multiplicative Chernoff bound (Appendix -A), to show that
$\mathbb{P}\left(E\geq\Psi_{{\mathcal{D}}_{\ell}}\right)$ $\hskip-0.75pt=\mathbb{P}(E\geq(1+\delta_{3})\mu_{E})$ (56)
$\hskip-0.75pt\leq\exp\left(-\big((1+\delta_{3})\log(1+\delta_{3})-\delta_{3}\big)\mu_{E}\right)$ (57)
$\hskip-0.75pt=\exp(-(1+\delta_{3})\log(1+\delta_{3})\mu_{E}(1-o(1)))$ (58)
$\hskip-0.75pt=\exp\left(-\Omega\Big(\ell\sqrt{\log n}\cdot\log\delta_{3}\Big)\right)$ (59)
$\hskip-0.75pt=\exp\left(-\Omega\left(\ell\sqrt{\log n}\cdot\log n\right)\right),$ (60)
@@
where (58) follows since $\delta_{3}=\omega(1)$, (59) follows since $(1+\delta_{3})\mu_{E}=\Psi_{{\mathcal{D}}_{\ell}}\geq\frac{1-\delta_{1}}{2}\ell\log n=\Omega(\ell\sqrt{\log n})$, and (60) follows since $\delta_{3}=\frac{n^{\Omega(1)}}{\sqrt{\log n}}$ and hence $\log\delta_{3}=\Omega(\log n)$. Lemma 7 then follows by recalling that $\mathbb{P}(T_{\exp}(\tilde{S})\leq T_{\exp}(\tilde{S}^{\prime}))=\mathbb{P}(E\geq\Psi_{{\mathcal{D}}_{\ell}})$ due to (42). ∎
@@
With this result established, we are now in a position to show that a union bound over all possible $({\mathcal{D}}_{\ell},{\mathcal{N}}_{\ell})$, followed by a second union bound over all the values of $\ell$, still leads to $o(1)$ probability of $\tilde{S}$ not explaining the most tests. This is formalized in the following lemma.
@@
Under the setup and notation of Lemma 7, [^5] In particular, we are again conditioning on (26) and (37), and the probability is over the randomness in the placements of non-defective items. for each $\ell\in\{1,2,\dots,5\eta k\}$, define
${\mathcal{E}}_{\ell}\coloneqq\bigcup_{({\mathcal{D}}_{\ell},{\mathcal{N}}_{\ell})}\left\{T_{\exp}(\tilde{S})\leq T_{\exp}((\tilde{S}\setminus{\mathcal{D}}_{\ell})\cup{\mathcal{N}}_{\ell})\right\},$ (61)
where each $({\mathcal{D}}_{\ell},{\mathcal{N}}_{\ell})$ pair implicitly satisfies $\lvert{\mathcal{D}}_{\ell}\rvert=\lvert{\mathcal{N}}_{\ell}\rvert$$\ell$, as the event that at least one way of swapping out $\ell$ items from $\tilde{S}$ with $\ell$ non-defective items leads to more tests being explained. Then, we have the following as $n\to\infty$:
$\mathbb{P}\left(\bigcup_{\ell=1}^{5\eta k}{\mathcal{E}}_{\ell}\right)\to 0.$ (62)
@@
We begin by deriving a bound on $\mathbb{P}({\mathcal{E}}_{\ell})$ for fixed $\ell$. For the event corresponding to any single choice of $({\mathcal{D}}_{\ell},{\mathcal{N}}_{\ell})$ in (61), the associated probability is upper bounded via Lemma 7. A simple counting argument reveals that there are ${(1-\eta)k\choose\ell}{n-k\choose\ell}$ such pairs, which can be further upper bounded by ${k\choose\ell}{n\choose\ell}$. Hence, the union bound gives
$\mathbb{P}({\mathcal{E}}_{\ell})$ $\leq{k\choose\ell}{n\choose\ell}\exp\left(-\Omega\big(\ell\sqrt{\log n}\cdot\log n\big)\right)$ (63)